OpenStudy (zzr0ck3r):

Just a question because its slow... We all love eulers Identity because it has 0, 1, e, pi ,and sqrt(-1) in it. What about 2? 2 is a pretty cool number also, and very important. so why not e^(2*pi*i) -1 = 0 ? (I just saw that video called "pi is wrong" and noticed this....and I think it would be a cooler tattoo)

OpenStudy (zzr0ck3r):

since were talking about it...

OpenStudy (anonymous):

Seconded.

OpenStudy (anonymous):

this is beautiful\[\large\color\green{\text{e}^{2\pi i}-1=0}\]

OpenStudy (anonymous):

2 is first prime number

OpenStudy (zzr0ck3r):

seems he missed the boat:)

OpenStudy (anonymous):

:)

OpenStudy (unklerhaukus):

i like the idea of \[\tau=2\pi\] but it would make more sense if the symbols were the other way around the symbol for pi looks like two taus joined together \[\qquad\qquad\qquad \pi\qquad\qquad\qquad\tau\tau\]

OpenStudy (zzr0ck3r):

ha good call lol